OpenStudy (anonymous):

solve for X: square root x^2-x-5=5 A. -5 B. -6 C. 6 D. -5,6

OpenStudy (anonymous):

the square root is over whole problem

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

do you mean \[\sqrt{x^2-x-5} = 5\]

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

so how would I go about solving it?

OpenStudy (anonymous):

I got -5

jimthompson5910 (jim_thompson5910):

\[\large \sqrt{x^2-x-5} = 5\] \[\large x^2-x-5 = 5^2\] \[\large x^2-x-5 = 25\] \[\large x^2-x-5 - 25 = 0\] \[\large x^2-x-30 = 0\]

jimthompson5910 (jim_thompson5910):

I'll let you finish

OpenStudy (anonymous):

yep I got my answer -5 that right?

jimthompson5910 (jim_thompson5910):

\[\large x^2-x-30 = 0\] \[\large (x+5)(x-6) = 0\]

OpenStudy (anonymous):

-5.6 is the answer?

jimthompson5910 (jim_thompson5910):

better, it's 6, -5 or -5, 6

OpenStudy (anonymous):

sorry haha that what I ment t type lol

jimthompson5910 (jim_thompson5910):

that's fine